If a cost-minimizing firm sells its services in an oligopolistic market, with a marginal product of capital of 10, a rental rate of capital of $5, and a wage rate of $20, what is the marginal product of labor?

1.4
2.20
3.40
4.60
5.80
1
Expert's answer
2021-08-09T11:07:02-0400

Answer:

The cost-minimizing condition is given below:

"\frac{MP_L}{w}=\frac{MP_K}{r}"


Where,

MPL=marginal product of labor

MPK=marginal product of capital

w=wage rate

r=rental rate of capital

Substitute the given values:


"\frac{MP_L}{20}=\frac{10}{5}"


"\frac{MP_L}{20}=2"


"{MP_L}={20}\times2=40"


The value of the marginal product of labor is 40. So, option 3 is correct.